How do I solve this NI question when MPC is increased by 20% and government spending is increased by 6%?

Given:

C= 200+0.6Y

I= 250

G= 200

R= 20


1
Expert's answer
2021-03-11T11:00:10-0500

Solution:

National Income (Y) = C + G + I + NX

C = 200 + 0.6Y

I = 250 – R = 250 – 20 = 230

G = 200 = has been increased by 6% = "200\times 1.06 = 212"

NEW G = 212

MPC = 0.6 = has been increased by 20% ="0.6\times 1.20 = 0.72"

NEW MPC = 0.72

R = 20

NI (Y) = 200 + 0.72Y + 212 + 230

NI (Y) = 200 + 212 + 230 + 0.72Y

NI (Y) = 642 + 0.72Y

Y – 0.72Y = 642

0.28Y = 642


Y = 2292.86

National Income = 2,292.86